-4(-3x - 2) + 3x = 9

Answer by algebrahouse.com(1659)

-4(-3x - 2) + 3x = 9
12x + 8 + 3x = 9 {used distributive property}
15x + 8 = 9 {combined like terms}
15x = 1 {subtracted 8 from each side}
x = 1/15 {divided each side by 15}

For more help from me, visit: www.algebrahouse.com